THE PURE ROTATIONAL SPECTRUM OF FeNC (X ^{6}Δ_{i}): IT HAD TO BE THERE!

\underline{M. A. FLORY} and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, Steward Observatory, University of Arizona, Tucson, AZ 85721.

The pure rotational spectrum of FeNC has been measured in its ground electronic and vibrational state using direct absorption methods. The radical was synthesized from gas-phase Fe(CO)_{5} and (CN)_{2} in an AC discharge and could also be produced using metal vapor and (CN)_{2} in a Broida oven. Eighteen rotational transitions have been recorded in the frequency range 150-505 GHz, many containing all six spin components. In addition, Renner-Teller effects have been observed for the excited vibrational v_{2} bending mode. The Ω = 3/2, 1/2, and -1/2 spin states appear to exhibit significant Λ-doubling. Preliminary results from isotopic substitution of ^{54}Fe, observed in natural abundance, indicate that the ground state structure is likely the linear isocyanide with a rotational constant near 4.1 GHz. Substitution of 13C is underway and will yield a definitive structure. The data are currently being analyzed with a case (a) Hamiltonian, and spectroscopic constants for rotation, fine structure, and Λ-doubling are being established.